OpenStudy (anonymous):

Estimate the radius of each circle with given circumference round to the nearest 10th 420in

OpenStudy (anonymous):

same problem. . .

OpenStudy (anonymous):

this one is 420in

OpenStudy (anonymous):

but, though, same process. .

OpenStudy (anonymous):

same idea... C=2*pi*r r=C/(2*pi) if you actually paid attention you'd be able to solve them quicker than you can type out the questions asking how to answer them

OpenStudy (anonymous):

Thanks